OpenStudy (anonymous):

Solve 5(2x + 1) = 12x – 1. You must show all work.

OpenStudy (blurbendy):

start by distributing the 5

OpenStudy (anonymous):

can u just show me step by step

OpenStudy (blurbendy):

5(2x + 1) = 12x – 1 10x + 5 = 12x - 1 Isolate x now by moving it to the right side of the equation (subtract 12 from both sides -2x + 5 = -1 Isolate x again by subtracting 5 from both sides -2x = -6 Isolate x again by dividing by -2 x = 3

OpenStudy (anonymous):

thank you

OpenStudy (blurbendy):

welcome